Calculate the pressure exerted by 0.0153 moles of CO2 gas in

a container of 1.467 L at 20.0 C.


1
Expert's answer
2021-03-12T05:59:55-0500

"P=\frac{mRT}{MV}"


R=8.31J mol/L

M= 44 g/mol=

V=2L

T=20 C=293.15 K

"m=M\times n=44\times0.0153=0.6732" g=0.0006732

"P=\frac{mRT}{MV}=\frac{6.732\times10^{-4}\times 8.31\times293.15}{44\times10^{-3}\times 1.467}=31.61" mPa